# It’s Giving Major Satisfaction: Why Fairness Matters for Developers

### v1

@@ IX Conclusion
Organizations should focus on applying all four fairness dimensions, particularly interpersonal fairness, within SE contexts like authorship, while tailoring practices to meet the needs of diverse demographic groups. Specific strategies have been outlined to address fairness concerns for software practitioners collectively (see Section VII-AVII-C) and based on their backgrounds (see Section VII-DVII-G). Future research could explore remedies for fairness problems that have already occurred, examine how fairness manifests in SE contexts, evaluate the effectiveness of suggested strategies, and bridge the gap between collective and demographic-specific fairness concerns.
@@
Our supplementary materials [62] contain source code used to retrieve and analyze data, results for all our RQs, and instructions how to replicate the results.
@@ Acknowledgments
The survey received ethics approval from the ethics committee at authors’ affiliations.

### v2

@@ 9. Conclusion
Organizations should focus on applying all four fairness dimensions, particularly interpersonal fairness, within SE contexts like authorship, while tailoring practices to meet the needs of diverse demographic groups. Specific strategies have been outlined to address fairness concerns for software practitioners collectively (see Section 7.17.3) and based on their backgrounds (see Section 7.47.7). Future research could explore remedies for fairness problems that have already occurred, examine how fairness manifests in SE contexts, evaluate the effectiveness of suggested strategies, and bridge the gap between collective and demographic-specific fairness concerns.
@@
Our supplementary materials (fig, 2024) contain source code used to retrieve and analyze data, results for all our RQs, and instructions how to replicate the results.
@@ Acknowledgments
The survey was approved by the ethics committee at the authors’ affiliations under ethics approval numbers 97126442 and UCL/CSREC/R/40.